OpenStudy (anonymous):

What is the approximate volume of a cone-shaped party hat a base whose circumference is 25.12 inches and height is 8 inches? a. 401.92 cubic inches b. 33.49 cubic inches c. 133.97 cubic inches d. 535.89 cubic inches

OpenStudy (anonymous):

Circumference = 2*pi*r Volume of Cone = 1/3 * pi * r^2 You're given the circumference; use it to find the radius. Plug that back into the volume formula for your answer.

OpenStudy (amistre64):

formula for any conic shape is: (1/3)(Area of Base)(height)

OpenStudy (anonymous):

so r=3.99 V=133,90 cubic inches ( C )

OpenStudy (anonymous):

Sorry, totally forgot to add *h to the volume formula!